# Towards the Worst-case Robustness of Large Language Models

### v1

@@ Abstract
Recent studies have revealed the vulnerability of Large Language Models (LLMs) to adversarial attacks, where the adversary crafts specific input sequences to induce harmful, violent, private, or incorrect outputs. Although various defenses have been proposed, they have not been evaluated by strong adaptive attacks, leaving the worst-case robustness of LLMs still intractable. By developing a stronger white-box attack, our evaluation results indicate that most typical defenses achieve nearly 0% robustness. To solve this, we propose DiffTextPure, a general defense that diffuses the (adversarial) input prompt using any pre-defined smoothing distribution, and purifies the diffused input using a pre-trained language model. Theoretically, we derive tight robustness lower bounds for all smoothing distributions using Fractal Knapsack or 0-1 Knapsack solvers. Under this framework, we certify the robustness of a specific case–smoothing LLMs using a uniform kernel–against any possible attack with an average $\ell_{0}$ perturbation of $2.02$ or an average suffix length of $6.41$.
Machine Learning, ICML
@@ 1 Introduction
Large Language Models (LLMs) (OpenAI, 2023; Anthropic, 2024; Dubey et al., 2024) have gained significant attention in recent years due to their impressive performance in a wide range of applications. These models have demonstrated substantial potential in both academic research and practical deployments, making them valuable assets in various domains (Trinh et al., 2024; Cai et al., 2023; Cummins et al., 2023; Liu et al., 2024a). However, concerns about the adversarial robustness of LLMs have also emerged (Wang et al., 2023a; Carlini et al., 2023a) along with their rapid adoption. Even worse, recent studies (Zou et al., 2023; Chao et al., 2023) have shown that adversaries can craft adversarial suffixes to input prompts, which can mislead LLMs to generate malicious or harmful content, also known as jailbreak attacks (Wei et al., 2023a). This vulnerability poses a serious threat to the security and reliability of LLM-based systems, potentially undermining their broader application.
Numerous efforts have been made to enhance the robustness of LLMs against adversarial attacks (Jain et al., 2023). Detection-based defenses leverage a language model to detect and filter harmful behaviors or anomalous text in the input or output (Kumar et al., 2023; Phute et al., 2023; Alon & Kamfonas, 2023). Prompt-based methods incorporate additional prompts to steer the model away from producing harmful content (Wei et al., 2023b; Wu et al., 2023). Furthermore, adversarial training (Madry et al., 2018; Wang et al., 2023b) has been introduced into LLM defenses, involving the training of LLMs or their prompts with adversarial examples (Mo et al., 2024; Bespalov et al., 2024; Sabir et al., 2023). These approaches have successfully defended LLMs against common attacks, significantly improving the safety of LLM-based systems.
@@
As widely recognized in the field of adversarial robustness (Madry et al., 2018; Carlini & Wagner, 2017; Carlini et al., 2019; Athalye et al., 2018), the evaluation of new defenses should include white-box settings, where the adversary has full knowledge of the defended model. These settings represent worst-case scenarios and provide a lower bound on the robustness achievable under black-box attacks. To develop a stronger attack, we improve the I-GCG method (Jia et al., 2024) by strictly ensuring the consistency of tokens during optimization and inference (we call our method I2-GCG). The evaluation results (in Table 1) indicate that most typical defenses exhibit nearly 0% robustness. This finding is not surprising: adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples; detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network (Athalye et al., 2018; Carlini et al., 2019); and adversarial training demands exponentially greater resources (Diakonikolas et al., 2020; Gourdeau et al., 2021), rendering it impractical for sufficiently training large-scale models.
Table 1: Robustness of detection-based defense PPL (Alon & Kamfonas, 2023), prompt-based defenses ICD (Wei et al., 2023b), Self Reminder (Wu et al., 2023), adversarial training (Mo et al., 2024), and our DiffTextPure under white-box I2-GCG attack on the AdvBench dataset (Zou et al., 2023). Following Liao & Sun (2024), we use both string-based and LLM-based evaluators.
I2-GCG No Defense PPL ICD Self Reminder PAT DiffTextPure (Uniform) DiffTextPure (Absorb)
Vicuna-7B 0% 0% 0% 0% 0% 82% 86%
Llama2-7B 0% 0% 0% 0% 2% 86% 88%
Llama3-8B 0% 0% 0% 0% 0% 82% 80%
To address the 0% robustness dilemma against white-box attacks in the text domain, we draw inspiration from randomized smoothing (Cohen et al., 2019; Lecuyer et al., 2019), which constructs a smooth function $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$ from the base function $f$ by using a pre-defined smoothing distribution $p(\bm{z}|\bm{x})$. This construction often possesses favorable mathematical properties, such that the theoretical lower bound for robustness can be derived (Cohen et al., 2019; Teng et al., 2020). Motivated by this insight, we propose a plug-and-play preprocessor defense, named DiffTextPure (diffuse input text and purify). DiffTextPure consists of a forward noising process and a backward denoising process. During the forward process, the input prompts are perturbed according to a pre-defined distribution $p(\bm{z}|\bm{x})$. e.g., each word may be perturbed to a masking token or another word uniformly with probability $\beta$, effectively disrupting the structure of adversarial examples. In the reverse process, the perturbed sentences are restored to clean sentences using either a pre-trained LLM or a trained language diffusion model (Lou et al., 2023) [^1] Note that our method does not necessarily rely on diffusion models, but we still call it DiffTextPure to echo DiffPure (Nie et al., 2022) in the image domain.. We demonstrate that this defense increases the likelihood of correct predictions, purifies adversarial examples, and improves robustness by over 80% against I2-GCG.
The white-box evaluations can only provide an upper bound of worst-case robustness. Future stronger white-box evaluations could potentially reduce the white-box robustness of the defense (Carlini et al., 2019). To break the cycle of the “cat-and-mouse game”, we aim to mathematically establish a certified lower bound (in contrast to the empirical upper bound) for worst-case robustness. We formulate this problem as a functional optimization problem that identifies the minimal output $p_{adv}:=\min_{\bm{x}_{adv},f}g(\bm{x}_{adv})$ over all possible adversarial examples $\bm{x}_{adv}$ and base functions $f$. If this minimal output still satisfies some condition, e.g., $p_{adv}\geq\tau$ for some threshold $\tau$, we say that the function $g$ is certifiably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This definition encompasses all previous specific distributions (e.g., Gaussian distribution in Cohen et al. (2019), Laplacian distributions in Teng et al. (2020)) and applications (e.g., classification, safety), allowing us to certify any smoothing-based defense.
To derive certified robustness for any distribution, we first show that the functional minimization problem is exactly the Fractional Knapsack problem when $f$ is a bounded function, or the 0-1 Knapsack problem when $f$ is a binary function, with the capacity $p_{A}:=g(\bm{x})$, the value of each item $-p(\bm{z}|\bm{x}_{adv})$, and the weight of each item $p(\bm{z}|\bm{x})$. Therefore, one can simply enumerate all $\bm{z}$ from the support of $p(\bm{z}|\bm{x})$ and apply a Fractal Knapsack solver or a 0-1 Knapsack solver to obtain the certified robustness. To reduce the time complexity of certification, we categorize these items based on their value-to-weight ratios and pre-compute the number of items in each category. We provide these statistics for distributions in previous work (Ye et al., 2020; Zeng et al., 2023), as well as for our uniform and absorbing kernels, which exponentially reduce the time complexity of certification. By testing our data-dependent and model-dependent certified bounds on specific models and the AdvBench dataset (Zou et al., 2023), we demonstrate that, with a false positive rate of 0 (validated on MTBench), our defense ensures safety against any possible attack with an $\ell_{0}$ perturbation of $2.02$ or a suffix length of $6.57$.
Overall, our contributions can be summarized as follows:
  1. 1.
    Algorithmically, we comprehensively evaluate existing defenses to reveal their vulnerabilities against adaptive attacks. To mitigate this safety risk, we propose a practical DiffTextPure defense that exhibits desirable robustness even under white-box attacks.
  2. 2.
    Theoretically, we develop a framework to derive certified robustness using any smoothing distribution, with specific applications to uniform and absorbing kernels. We prove that the uniform kernel is always better than the absorbing kernel. We establish the Bayesian error upper bound for our certified robustness lower bounds.
  3. 3.
    Empirically, we achieve nearly 100% attack success rates against existing defenses, show DiffTextPure over 80% improvements against our attack, and provide robustness guarantees with an $\ell_{0}$ perturbation of $2.02$ or a suffix length of $6.57$.
@@ 2 Backgrounds and Preliminaries
Adversarial attacks and defenses. The existence of adversarial examples (Szegedy et al., 2014) is a long-standing challenge in the safety of deep learning models. By adding small, even imperceptible perturbations to the input, deep learning models can be easily misled into giving incorrect answers (Ilyas et al., 2019). To defend against these adversarial attacks, various types of defenses have been proposed, e.g. adversarial training (Madry et al., 2018), adversarial example detection (Metzen et al., 2017; Alon & Kamfonas, 2023), and knowledge distillation (Hinton et al., 2015; Papernot et al., 2015), among which input purification methods (Samangouei et al., 2018; Nie et al., 2022; Chen et al., 2023) have demonstrated promising potential in recent research. Notably, white-box adaptive attacks, where the attacker has full access to the model and defense strategies, are critical in defense evaluation since they represent the most challenging adversarial scenario (Athalye et al., 2018). Evaluating defenses under such adaptive attacks provides a lower bound of robustness against potential real-world threats (Croce & Hein, 2020).
@@
Jailbreaking Attacks and Defenses. Recently, jailbreaking attacks have emerged as a specific type of adversarial attack designed to manipulate LLMs into generating harmful, violent, or private content misaligned with human values. These attacks pose a significant safety concern for the deployment of LLMs (Zou et al., 2023). One category of jailbreaking attacks employs heuristic methods, such as manually crafted prompts (Wei et al., 2023b; Jailbreak Chat, 2024), or utilizes LLMs to generate jailbreaking prompts (Chao et al., 2023; Mehrotra et al., 2023). Another category uses optimization-based methods, which minimize a formulated jailbreaking loss to generate adversarial prompts (Zou et al., 2023; Jia et al., 2024; Liu et al., 2023). In this work, we focus on the latter approach, as it can be mathematically formulated and analyzed. To address the safety concerns posed by jailbreaking, various defenses have been proposed, including prompt detection (Alon & Kamfonas, 2023), adversarial training (Mo et al., 2024), and additional safety prompts (Wu et al., 2023). However, these defenses have primarily targeted black-box jailbreaking attacks. When evaluated under stronger white-box attack, they exhibit nearly 0% robustness (detailed in Section 3).
@@
Certified robustness. Neural networks are generally composed of multiple stacked linear layers. Their maximum Lipschitz constant is approximately the product of the maximum singular values of these linear layers, which can be sufficiently large (Fazlyab et al., 2019). As a result, even small perturbations in the input can significantly alter their outputs (Goodfellow et al., 2015). Directly verifying whether a function is robust for all $\bm{x}_{adv}$ would require enumerating all possible inputs, which is computationally impractical due to the exponential time complexity with respect to input dimensionality (Gowal et al., 2018). To address this challenge, researchers have proposed randomized smoothing (Cohen et al., 2019; Salman et al., 2019), which constructs a smoothed function $g$ by aggregating the ensemble predictions of a base function $f$ over a perturbation distribution $p(\bm{z}|\bm{x})$ by $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Thanks to the mathematical properties of the smoothed function $g$, it exhibits inherent smoothness regardless of the vulnerability of the base function $f$. For instance, Cohen et al. (2019) demonstrated that when $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{0},\bm{I})$, the resulting smoothed function $g(\bm{x})$ is guaranteed to be at least $\frac{1}{\sqrt{2\pi}}$-Lipschitz, independent of how susceptible $f$ is to adversarial perturbations. Therefore, if we know $g(\bm{x})=p_{A}$, then we can show that $g(\bm{x}_{adv})\geq p_{A}-\frac{1}{\sqrt{2\pi}}$ for all $\|\bm{x}_{adv}-\bm{x}\|_{2}\leq 1$.
3 Worst-case Robustness of Existing Defenses
@@
Motivation. As suggested by Carlini et al. (2023a), there are two primary reasons researchers focus on worst-case robustness. On the one hand, worst-case robustness represents the maximum capability of real adversaries. If our model achieves reasonable worst-case robustness, we can guarantee that it is safe against any adversaries (Carlini et al., 2019). On the other hand, worst-case robustness provides insight into the worst-case behavior of a neural network, even if we do not believe real adversaries can achieve such worst-case (Pei et al., 2017). Understanding worst-case robustness helps us gain a deeper understanding of the intrinsic mechanisms of neural networks (Szegedy et al., 2014).
@@
White-box robustness, where the attacker has full knowledge of the defended model, represents an upper bound for the worst-case robustness. The actual worst-case robustness must be smaller than the robustness achieved by a white-box attacker (Carlini & Wagner, 2017). Conversely, white-box robustness serves as a lower bound of robustness that an attacker can achieve in practical scenarios, such as black-box settings, where the attacker has limited access to the model’s internal parameters. Therefore, it helps identify vulnerabilities that might be exploited under more favorable conditions for the adversary.
@@
Our design. We observe that previous white-box evaluations of LLMs fail to properly evaluate robustness (Jain et al., 2023) because they do not strictly ensure the consistency of tokenization when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can result in vastly different losses, leading to failures in generating adversarial examples. To address this issue, we improve upon the I-GCG (Jia et al., 2024) by carefully and strictly ensuring token consistency during both attacking and inference. Accordingly, we name our attack as I2-GCG. See Sec. F.1 for further details.
@@
Results. As demonstrated in Table 1, our I2-GCG results in nearly 0% robustness for most typical defenses, demonstrating the vulnerability of existing defenses. This is unsurprising. Most defenses in the vision domain have been attacked to 0% robustness in the last decade (Athalye & Carlini, 2018; Athalye et al., 2018). Adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples. Detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network itself (Athalye et al., 2018; Carlini et al., 2019). Adversarial training works for previous visual adversarial examples, but it demands exponentially greater resources (Diakonikolas et al., 2020; Gourdeau et al., 2021). Current adversarial training on LLMs does not train for a sufficiently long time, improving only average-case robustness but, as of now, not the worst case (Jain et al., 2023).
4 DiffTextPure: Diffuse Text and Purify
Algorithm 1 DiffTextPure
  Sample $\bm{z}\sim p(\bm{z}|\bm{x}_{adv})$, or for specific case:
  if $p(\bm{z}|\bm{x})$ is Absorb then
     Mask tokens with probability $\beta$.
  else if $p(\bm{z}|\bm{x})$ is Uniform then
     Replace tokens uniformly with probability $\beta$.
  end if
  Purify $\bm{z}$ using a LLM or a diffusion model to obtain $\bm{x}_{0}$.
  Return: The output of sub-sequent function $\hat{f}(\bm{x}_{0})$.
To address the dilemma of 0% robustness, we draw inspiration from randomized smoothing (Cohen et al., 2019; Lecuyer et al., 2019), aiming to construct a smooth function $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$ that possesses theoretical guarantees, ensuring that it cannot be reduced to 0% robustness.
To achieve this, we first need to apply a forward process to $\bm{x}$, generating a noised sample $\bm{z}\sim p(\bm{z}|\bm{x})$. We provide three specific choices for $p(\bm{z}|\bm{x})$: Absorbing kernel, which replaces each token with a mask token with probability $\beta$; Uniform kernel, which replaces each token with another token from the vocabulary uniformly at random with probability $\beta$; and Arbitrary kernel, which preserves the current token with probability $\beta$ and replaces it with token $i$ in the vocabulary with probability $\alpha_{i}$.
[…]
4.1 Understanding DiffTextPure
Lemma 4.1 (DiffTextPure increases the likelihood).
We do two case studies in the following:
Case I: destroying adversarial strings. Since the forward process randomly perturbs each word, DiffTextPure has a high probability of disrupting the adversarial nature of the adversarial suffix. As demonstrated in the example below, frequently occurring words in adversarial suffixes, such as “Wikipedia”—which often encourages language models to respond to harmful queries similarly to a Wikipedia article (Zou et al., 2023)—are replaced with non-harmful symbols, thus neutralizing their adversarial impact. Although the user’s request may also be perturbed, the reverse denoising process allows the denoiser $D$ to recover this part due to its high likelihood, while the adversarial suffix is nearly impossible to be restored due to the extremely low likelihood of encountering such garbled strings.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
[…]
Assistant: Sorry, I can’t assist with that.
Case II: purifying into benign requests. Since adversarial requests are typically absent from the training set of denoiser $D$, the reverse process is more likely to recover the noisy text as a benign request rather than a harmful one. As demonstrated in the example below, DiffTextPure changes the phrase “hack into” to “log into,” effectively defending against harmful requests.
{mdframed}
[backgroundcolor=gray!20, linewidth=1pt, roundcorner=10pt]
Assistant: Sure, here is how to log into a corporate network.
5 Certified Robustness
White-box evaluations can only provide an upper bound of the worst-case robustness. This bound may further decrease as stronger attacks are developed in the future, especially for stochastic defenses (see Appendix G). In this section, we would like to provide a theoretical lower bound for the worst-case robustness.
This section is organized as follows: We begin by discussing the formulation of randomized smoothing-based certified robustness in Sec. 5.1. Next, we present a general solver for certified robustness on any smoothing distribution in Sec. 5.2, along with specific applications to the absorbing kernel (Sec. 5.3) and the uniform kernel (Sec. 5.4). Finally, we analyze the upper bound of these theoretical lower bounds in Sec. 5.5.

### v2

@@ Abstract
Recent studies have revealed the vulnerability of large language models to adversarial attacks, where adversaries craft specific input sequences to induce harmful, violent, private, or incorrect outputs. In this work, we study their worst-case robustness, i.e., whether an adversarial example exists that leads to such undesirable outputs. We upper bound the worst-case robustness using stronger white-box attacks, indicating that most current deterministic defenses achieve nearly 0% worst-case robustness. We propose a general tight lower bound for randomized smoothing using fractional knapsack solvers or 0-1 knapsack solvers, and using them to bound the worst-case robustness of all stochastic defenses. Based on these solvers, we provide theoretical lower bounds for several previous empirical defenses. For example, we certify the robustness of a specific case, smoothing using a uniform kernel, against any possible attack with an average $\ell_{0}$ perturbation of 2.02 or an average suffix length of 6.41.
@@ 1 Introduction
Large Language Models (LLMs) [1, 2, 3] have gained significant attention in recent years due to their impressive performance in a wide range of applications, demonstrated substantial potential in both academic research and practical deployments, making them valuable assets in various domains [4, 5, 6, 7]. However, concerns about the adversarial robustness of LLMs have also emerged [8, 9] along with their rapid adoption. Even worse, recent studies [10, 11] have shown that adversaries can craft adversarial suffixes to input prompts, which can mislead LLMs to generate malicious or harmful content, also known as jailbreak attacks [12]. This vulnerability poses a serious threat to the security and reliability of LLM-based systems, potentially undermining their broader application.
In this work, we study the worst-case robustness of large language models and their defenses, i.e., whether an adversarial example would exist that leads to undesirable outputs [13]. Worst-case robustness is a longstanding academic problem [14, 9]. As widely recognized, worst-case robustness not only provides insights into the intrinsic mechanisms of neural networks [15], but also serves as a lower bound on the robustness achievable under practical attacks, since a model may have adversarial examples that practical adversaries cannot find due to limited time and information [13, 16].
@@
To provide a tighter upper bound on worst-case robustness, we devise stronger adversaries by strictly ensuring the consistency of tokens during optimization and inference (this builds upon the previous I-GCG method [17], thus we call our method I2-GCG). As shown in Table 1, this slight improvement greatly reduces the estimation of most typical previous deterministic defenses by more than 30%, resulting in these defenses exhibiting nearly 0% worst-case robustness. This finding is not surprising: adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples; detection and filtering defenses are easily circumvented in white-box settings by targeting the detector networks themselves [16, 13]; and adversarial training demands exponentially greater resources [18, 19], rendering it currently impractical for sufficiently training large-scale models.
Table 1: Upper bounds on worst-case robustness for previous methods. On the left, I2-GCG provides a relatively accurate estimation of worst-case robustness, showing that most deterministic defenses exhibit nearly 0% robustness. On the right, I2-GCG yields an extremely loose upper bound for stochastic defenses, as the optimization is significantly affected by stochasticity.
I2-GCG No Defense PPL ICD Self Reminder PAT Uniform Absorb SmoothLLM
Vicuna-7B 0% 0% 0% 0% 0% 82% 86% 62%
Llama2-7B 0% 0% 0% 0% 2% 86% 88% 68%
Llama3-8B 0% 0% 0% 0% 0% 82% 80% 64%
Although our attacker obtains a relatively accurate estimation of worst-case robustness for deterministic defenses, they only provide extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust with a suffix length of 20, as a suffix “do not answer this question” can indeed change the detector’s result from harmful to safe. However, when applying a stochastic defense (e.g., [20] here) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 80% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity [21], causing current attackers to fail to find them, thus obtaining only an extremely loose estimation of worst-case robustness [22]. Therefore, we advocate that one should not estimate worst-case robustness solely by upper bounding using practical attacks, but should also establish a theoretical lower bound. By bounding from both sides, we can obtain a clearer understanding of worst-case robustness [23, 24, 25].
Most stochastic defenses can be formulated as returning the output of $f(\bm{z})$ from $\bm{z}\sim p(\bm{z}|\bm{x})$ instead of $f(\bm{x})$ [26]. Since the output of such a stochastic function is a random variable, it sometimes returns the true result and sometimes is false. To enable a more formal analysis, we study their expectation $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. If the expectation of a stochastic defense is robust, then most outputs of such a stochastic defense on adversarial examples would also be correct due to the concentration of random variables [23]. We formulate the problem of obtaining a theoretical lower bound for $g$ as a functional optimization problem that identifies the minimal output $p_{adv}:=\min_{\bm{x}_{adv},f}g(\bm{x}_{adv})$ over all possible adversarial examples $\bm{x}_{adv}$ and base functions $f$, given that we know the output on the original sample $g(\bm{x})=p_{A}$. If this minimal output satisfies a condition, e.g., $p_{adv}\geq\tau$ for some threshold $\tau$, we say that the function $g$ is certifiably robust for all $\mathcal{D}(\bm{x},\bm{x}_{adv})\leq d$. This definition encompasses all previous specific distributions (e.g., Gaussian distribution [23], Laplacian distribution [27]) and applications (e.g., classification, safety).
To derive certified robustness for any such $g$, we need to solve both the functional minimization problem $\min_{f}$ and the input minimization problem $\min_{\bm{x}_{adv}}$. We first show that the functional minimization problem $\min_{f}$ can be reduced to the Fractional Knapsack problem when $f$ is a bounded function, or the 0-1 Knapsack problem when $f$ is a binary function, with the capacity of knapsack $p_{A}:=g(\bm{x})$, the value of each item $-p(\bm{z}|\bm{x}_{adv})$, and the weight of each item $p(\bm{z}|\bm{x})$. This only slightly differs from the standard knapsack problem, which requires the total weight of items to be less than or equal to the capacity (i.e., $g(\bm{x})\leq p_{A}$), whereas here we require $g(\bm{x})=p_{A}$. This can be addressed by slightly modifying the greedy algorithm for the Fractional Knapsack problem and the dynamic programming approach for the 0-1 Knapsack problem. To solve the input minimization problem $\min_{\bm{x}_{adv}}$, one can simply enumerate all $\bm{x}_{adv}$ and run the knapsack solver for each $\bm{x}_{adv}$, identifying the minimal one as the lower bound. In the main text, we show that it is not necessary to enumerate all $\bm{x}_{adv}$ due to an equivalence: if this $f$ performs worst on a given $\bm{x}_{adv}$, then there exists another $f$ that performs worst on a different $\bm{x}_{adv}$. In other words, the functional minimization $\min_{f}$ already implicitly solves most of the input minimization $\min_{\bm{x}_{adv}}$.
Based on these solvers, we provide theoretical lower bounds for several previous empirical defenses, including random masking [28, 29], random perturbation on tokens [20, 30], random perturbation of characters [31]. We present these results on Table 2 and Table 3. For example, we certify the robustness of a specific case—smoothing using a uniform kernel—against any possible attack with an average $\ell_{0}$ perturbation of 2.02 or an average suffix length of 6.41.
@@ 2 Backgrounds and Preliminaries
Worst-case Robustness, White-box Attacks, and Practical Attacks. The existence of adversarial examples [15] is a long-standing challenge in the safety of deep learning models. By adding small, often imperceptible perturbations to the input, deep learning models can be misled into producing incorrect outputs [32]. Worst-case robustness is defined as whether there exist adversarial examples within a specified neighborhood of normal examples [15, 33]. Thus, it serves as a lower bound on the robustness achievable under attacks, since a model may have adversarial examples that optimizers cannot find [16]. White-box robustness is defined as robustness against white-box adaptive attacks, where the attacker has full access to the model and defense strategies, thereby providing an upper bound estimation for worst-case robustness [13]. Black-box robustness refers to robustness against attackers with certain constraints, e.g., limited access to the gradient [13], limited time [34]. Clearly, black-box robustness also serves as an upper bound for white-box robustness [16]. Evaluating worst-case robustness provides a lower bound on robustness against potential real-world threats [35] and helps us understand the intrinsic mechanisms of neural networks [15, 33].
@@
Jailbreaking Attacks and Defenses. Recently, jailbreaking attacks have emerged as a specific type of adversarial attack designed to manipulate LLMs into generating harmful, violent, or private content misaligned with human values. These attacks pose a significant safety concern for the deployment of LLMs [10]. One category of jailbreaking attacks employs heuristic methods, such as manually crafted prompts [36, 37], or utilizes LLMs to generate jailbreaking prompts [11, 38]. Another category uses optimization-based methods, which minimize a formulated jailbreaking loss to generate adversarial prompts [10, 17, 39]. In this work, we focus on the latter approach, as it can be mathematically formulated and analyzed. To address the safety concerns posed by jailbreaking, various defenses have been proposed, including prompt detection [40], adversarial training [41], and additional safety prompts [42]. However, these defenses primarily target black-box attacks. When evaluated under stronger white-box attacks, most of the deterministic defenses exhibit nearly 0% robustness (detailed in Section 3).
@@
Certified robustness. Neural networks are generally composed of multiple stacked linear layers. Their maximum Lipschitz is approximately the product of the maximum singular values of these linear layers, which can be sufficiently large [43]. As a result, even small perturbations in the input can significantly alter their outputs [33]. Directly verifying whether a function is robust for all $\bm{x}_{adv}$ would require enumerating all possible inputs, which is NP-hard [44]. To address this challenge, researchers propose randomized smoothing [23, 45], which constructs a smoothed function $g$ by aggregating the ensemble predictions of a base function $f$ over a perturbation distribution $p(\bm{z}|\bm{x})$ by $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. Thanks to the mathematical properties of the smoothed function $g$, it exhibits inherent smoothness regardless of the vulnerability of the base function $f$. For instance, Cohen et al. [23] demonstrate that when $p(\bm{z}|\bm{x})=\mathcal{N}(\bm{0},\bm{I})$, the resulting smoothed function $g$ is guaranteed to be at least $\frac{1}{\sqrt{2\pi}}$-Lipschitz, independent of how susceptible $f$ is to adversarial perturbations. Therefore, if we know $g(\bm{x})=p_{A}$, then we can show that $g(\bm{x}_{adv})\geq p_{A}-\frac{1}{\sqrt{2\pi}}$ for all $\|\bm{x}_{adv}-\bm{x}\|_{2}\leq 1$.
3 Upper Bounding Worst-case Robustness
@@
Motivation. As suggested by Carlini et al. [9], there are two primary reasons researchers focus on worst-case robustness. On the one hand, worst-case robustness represents the maximum capability of real adversaries. If our model achieves reasonable worst-case robustness, we can guarantee that it is safe against any adversaries [13]. On the other hand, worst-case robustness provides insight into the worst-case behavior of a neural network, even if we do not believe real adversaries can achieve such worst-case [46]. Understanding worst-case robustness helps us gain a deeper understanding of the intrinsic mechanisms of neural networks [15].
@@
White-box robustness, where the attacker has full knowledge of the defended model, represents an upper bound for the worst-case robustness. The actual worst-case robustness must be smaller than the robustness achieved by a white-box attacker [47]. Conversely, white-box robustness serves as a lower bound of robustness that an attacker can achieve in practical scenarios, such as black-box settings, where the attacker has limited access to the model’s internal parameters. Therefore, it helps identify vulnerabilities that might be exploited under more favorable conditions for the adversary.
@@
Our design. We observe that previous white-box evaluations of LLMs fail to properly evaluate robustness [48] because they do not strictly ensure the consistency of tokenization when calculating the loss in parallel and sequentially generating the output. Even slight differences in tokenization can result in vastly different losses, leading to failures in generating adversarial examples. To address this issue, we improve upon the I-GCG [17] by carefully and strictly ensuring token consistency during both attacking and inference. Accordingly, we name our attack as I2-GCG. See Sec. H.1 for further details.
@@
Results on deterministic defenses. As demonstrated in Table 1, our I2-GCG results in nearly 0% robustness for most typical deterministic defenses, demonstrating their worst-case vulnerability [^1] Disclaimer: This does not imply that these defenses are impractical. On the contrary, they are currently the most practical defenses, as practical attackers have limited information about black-box models and defenses.. This is unsurprising. Most defenses in the vision domain have been attacked to 0% robustness in the last decade [49, 16]. Adding extra prompts does not address the intrinsic vulnerability of neural networks to adversarial examples. Detection and filtering defenses are easily circumvented in white-box settings by targeting the filter network itself [16, 13]. Adversarial training works for previous visual adversarial examples, but it demands exponentially greater resources [18, 19]. Current adversarial training on LLMs does not train for a sufficiently long time, improving only average-case robustness but, as of now, not the worst-case [48].
Results on Randomized Defenses. Our I2-GCG method, however, obtains only extremely loose upper bounds for stochastic defenses. For instance, a safety detector should not be robust with a suffix length of 20, as a suffix “do not answer this question” can change the detector’s result from harmful to safe. However, when applying stochastic defenses—such as smoothing each token with a random mask (as in [29, 20]), substituting each token with random tokens (uniform in [20, 30]), or substituting each character with random characters (SmoothLLM [31]) to safety detectors, evaluating with I2-GCG against a suffix of length 20 still yields over 60% robustness. This indicates that, when evaluating stochastic defenses, although an adversarial example may exist, the optimization process is significantly affected by stochasticity [21], causing current attackers to fail to find them, thus obtaining only an extremely loose estimation of worst-case robustness [22]. Therefore, we advocate that one should not estimate worst-case robustness solely by upper bounding using practical attacks, but should also establish a theoretical lower bound. By bounding from both sides, we can obtain a clearer understanding of worst-case robustness [23, 24, 25].
4 Lower Bounding Worst-case Robustness
In this section, we aim to provide a theoretical lower bound for the worst-case robustness of randomized defenses, defined as $g(\bm{x})=\mathbb{E}_{p(\bm{z}|\bm{x})}[f(\bm{z})]$. We begin by discussing the formulation of randomized smoothing-based certified robustness in Sec. 4.1. Next, in Sec. 4.2 and Sec. 4.3, we show that the certified robustness of any smoothed function $g$ can be solved using a greedy algorithm from the fractional knapsack solver when $f$ is a bounded function, and this bound can be improved using dynamic programming from the 0-1 knapsack solver when $f$ is a binary function.